SLOW ELECTRON VELOCITY-MAP IMAGING OF La_{2}(C_{6}H_{6}) AND La(C_{6}H_{6})_{2}

\underline{RUCHIRA SILVA} AND DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Slow electron velocity-map imaging (SEVI) was used to study the structures and electronic states of La_{2}(C_{6}H_{6}) and La(C_{6}H_{6})_{2} complexes formed in a metal cluster beam source. Electron spectra obtained from SEVI have the energy resolution of ΔE/eKE ≈ 2% at eKE=400cm^{−1}. The SEVI technique offers much higher data collection efficiency than pulsed-field ionization zero electron kinetic energy spectroscopy. From the SEVI spectra, adiabatic ionization energies were measured to be 32141(5) cm^{−1} for La_{2}(C_{6}H_{6}) and 39033(5) cm^{−1} for La(C_{6}H_{6})_{2}. The most active vibrational transition for both complexes was identified to be a metal-ligand stretching mode with a frequency of 180 cm^{−1} in the ion state. In addition, a benzene ring out-of-plane deformation mode was measured to be 303 cm^{−1} for [La_{2}(C_{6}H_{6})]^{+} and 408 cm^{−1} for [La(C_{6}H_{6})_{2}]^{+}. By combining the spectra with theoretical calculations, we identified the ^{2}A_{g}←^{1}A_{g} transition of La_{2}(C_{6}H_{6}) (C_{2h}) and ^{3}A←^{2}A of La(C_{6}H_{6})_{2} (C_{1}). By measuring the anisotropy parameter (β) from photoelectron angular distribution, we found that the outgoing electron was from a largely La 6s-based molecular orbital in both complexes.